Student A prepared a 0.038M of the stock sodium salicylate solution using a pipette, transferred 5mL of the 0.038M sodium salicylate stock solution into a 100mL volumetric flask labelled as sample 1. The solution was diluted to the 100ml mark with acidified 0.02M iron (III) chloride solution. What is the concentration of Iron (III) salicylate complex (in M) in the sample 1 flask. Enter the answer in 4 decimal places
1
Expert's answer
2020-09-29T06:56:04-0400

STEP I: BALANCED EQUATION


3C7H5NaO3(aq) + FeCl3(aq) "\to" Fe(C7H5O3)3 (aq) + 3NaCl(aq)


From the equation, 3 moles of C7H5NaO3 is required to react with only 1 mole of FeCl3 to produce 1 mole of the Iron (III) salicylate complex.


STEP II : CALCULATE MOLES OF C7H5NaO3


From the concentration C7H5NaO3 ;

0.038 moles were in 1000mL

how many moles would be in 5mL?


"=\dfrac{0.038molx5mL}{1000mL}=1.9x10^-4moles"


STEP III : CALCULATE MOLES OF FeCl3 REQUIRED


From mole ratio of 1:3

Moles of FeCl3 required"=\dfrac{1}{3}x1.9x10^-4mol =6.33x10^-5moles"


But the moles of FeCl3 already available in the flask are?

From the concentration and volume of FeCl3 we know that; (95 mL of FeCl3 was added)

0.02 moles are in 1000mL

How many moles therefore would be in the 95mL?

"=\dfrac{0.02molx95mL}{1000mL}=1.9x10^-3moles"


That means that FeCl3 was in excess and the mass of the final moles of the product will depend on the moles of the sodium salicylate solution (the limiting reagent).


STEP IV : CALCULATE MOLES OF THE PRODUCT

Moles of Iron (III) salicylate complex from the 1:1 mole ratio is therefore:


"=\dfrac{1}{1}xmoles ofC7H5NaO3"


"=\dfrac{1}{1}x1.9x10^-4"


= 1.9 x 10-4 moles.


STEP V : CONVERT THESE MOLES TO MOLARITY


But these moles (1.9x10-4) were in 100mL of solution?

How many moles would be in 1000mL

"=\dfrac{(1.9x10^-4mol)x1000mL}{100mL}=1.9x10^-3M"



The concentration of the Iron (III) salicylate complex is therefore; 1.9x10-3M